# Incremental Learning for Robot Shared Autonomy

### v1

@@ III-C Robot Action Generation Model
Our robot action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$, shown in Fig. 2, produces two types of actions at each time step: an intermediate action $a_{m,t}=\mathcal{F^{\prime}}_{\theta^{\prime}}(s_{t})$ without user guidance and a final action $a_{f,t}=\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$ with user guidance. Here, $\mathcal{F^{\prime}}_{\theta^{\prime}}$ and ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$ represent different subsets of the overall model $\mathcal{F}_{\theta}$, as illustrated in Fig. 2. While only the action with user guidance is executed, having an intermediate output is crucial during the incremental learning phase, as we will discuss in Section III-D.
We pretrain $\mathcal{F}_{\theta}$ in an end-to-end fashion with synthetic kinematic trajectories (Section III-B). We achieve the balance between task completion and the user’s sense of control through a loss function defined as $\mathcal{L}=\alpha\mathcal{L}_{demo,m}+\beta\mathcal{L}_{demo,f}+\gamma\mathcal{L}_{direc}+\delta\mathcal{L}_{order}$, where $\alpha$, $\beta$, $\gamma$ and $\delta$ are weights that adjust the contribution of each loss term. We set $\alpha=\beta=1$ and $\gamma=\delta=100$.
@@
where $a_{r,t}$ is the robot action in the kinematic trajectories. These two loss terms supervise the intermediate and final outputs respectively, guiding the robot to take actions towards task completion. To further ensure that the final executed action aligns with the user’s input, we introduce two additional loss terms, action direction loss $\mathcal{L}_{direc}$ and action ordering loss $\mathcal{L}_{order}$, to supervise $a_{f,t}$.
@@
III-D Incremental Learning
ILSA starts with the robot action generation model $\mathcal{F}_{\theta}$ pretrained on kinematic trajectories. Our goal is to incrementally finetune $\mathcal{F}_{\theta}$ using interaction data collected during deployment.
During deployment, at each time step, we obtain the user action $a_{u,t}$ from the control interface and input it into $\mathcal{F}_{\theta}$ to generate the final robot action $a_{f,t}$. Since the pretrained model may produce errors, we compute the cosine similarity between $a_{u,t}$ and $a_{f,t}$. If the similarity is above an empirically set threshold (0.5), we execute $a_{f,t}$. If it falls below the threshold, indicating a discrepancy, we alert the user by briefly pausing the robot’s motion. If the user persists with the same input, the robot will execute $a_{u,t}$ instead of $a_{f,t}$.
As the robot continues to operate using this process, we collect interaction data to finetune $\mathcal{F}_{\theta}$. However, simply updating the entire model with these raw trajectories has limitations. In the following sections, we discuss these challenges and propose corresponding solutions.
III-D1 Corrected Trajectory Generation
Using raw trajectories of user interactions to update $\mathcal{F}_{\theta}$ can be suboptimal, as our experiments show that when users overwrite $a_{f,t}$ with $a_{u,t}$, it is typically to correct errors previously made by the system.
For example, in the case shown in Fig. 1(a), when the robot arm collides with a shelf while retrieving a cup containing cereal, the user wants to command the robot to move down to avoid further mistakes. This action is an outlier unexpected by the robot, hence the user needs to overwrite $a_{f,t}$ with $a_{u,t}$. This process is only necessary because previously, the pretrained model didn’t account for environmental collisions and mistakenly attempted a linear trajectory to move next to the bowl, forcing the user to correct it.
As illustrated in this example, raw trajectories contain both policy errors and user corrections that wouldn’t be necessary without the errors. Directly using these trajectories for training risks misleading the model into treating these corrective actions as optimal behavior to be replicated.
To address this, we generate a “corrected trajectory” for each raw trajectory observed during human use. As shown in Fig. 3, when the user overwrites $a_{f,t}$ (red segments in Fig. 3), let $t_{1}$ be the start of the overwrite and $t_{2}$ the time when the robot resumes executing $a_{f,t}$. Before $t_{1}$, if the robot executes $a_{f,t}$ for a period (blue segments), let $t_{0}$ be the start of that period. Let $p_{t_{0}}$, $p_{t_{1}}$, and $p_{t_{2}}$ represent the robot’s positions at these times.
III-D2 Layered Supervision
After generating the corrected trajectory, a straightforward approach is to update the model using only the newly generated trajectory. However, for rapid learning, ILSA is designed to be updated after each new interaction, and training on only one or a few new trajectories can result in overfitting, gradient noise, and catastrophic forgetting [11, 26, 18].
Incorporating synthetic kinematic trajectories into the training set can help address this issue. Instead of directly combining the kinematic trajectories with new trajectories, which may lead to data imbalance, we propose a layered supervision approach:
  1. 1.
    Final Output Supervision: We only use the new user-generated trajectories to supervise $a_{f,t}$, i.e., finetune ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$.
  2. 2.
    Intermediate Output Supervision: Both the synthetic kinematics trajectories and new user-generated trajectories are used to supervise $a_{m,t}$, i.e., finetune $\mathcal{F^{\prime}}_{\theta^{\prime}}$.
This approach is illustrated by the red arrows in Fig. 2. It integrates the benefits of both synthetic kinematic trajectories and new trajectories, while prioritizing the model’s adaptation to the newly generated user data during fine-tuning. Compared to other commonly used techniques in human-in-the loop machine learning methods such as assigning different weights to different types of data [25, 22, 21], ILSA avoids the need for data reweighting or dataset subsampling.
III-D3 Partial Model Update
To improve the robustness of the finetuning phase, we update only the transformer in the robot action generation model, which maps the task state $s_{t}$ into state features, while freezing other parts of the architecture. The rationale is that the initial model’s inefficiencies mainly stem from the simplistic linear nature of the synthetic kinematic trajectories, which fail to take full consideration of all environmental factors. Updating the transformer, responsible for modeling environmental variations, can sufficiently address this limitation.
IV Experiments
IV-A Tasks and Experiment Settings
All experiments are conducted with a 6-DOF Kinova Gen3 robotic arm, and users use an Xbox controller to send commands to the robot. We test the effectiveness of our model on two complex, multi-stage tasks:
IV-A1 Cereal Pouring
IV-A2 Pill Bottle Storage
Open a drawer, pick up a pill bottle behind the drawer, and place it back inside the opened drawer.
IV-B Ablation Study
Prior to the user study, we conducted an ablation study on the cereal pouring task to verify the design choices of ILSA. The ablated components and alternative designs evaluated are:
  • ILSA finetunes on corrected trajectories: A. Finetune on original uncorrected trajectories.
  • ILSA uses new data to supervise both $\bm{a_{f,t}}$ and $\bm{a_{m,t}}$, and kinematic trajectories to supervise only $\bm{a_{m,t}}$: B1. Use new data only to supervise both $a_{f,t}$ and $a_{m,t}$; B2. Use both new data and kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$; B3. Use new data and randomly selected 10 kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$; B4. Use both new data and kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$, reweighted with $w_{new}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{new}|}$ and $w_{kinematic}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{kinematic}|}$, where $|D_{new}|$ and $|D_{kinematic}|$ represent the number of data points in the new and kinematic trajectories, respectively (This design is inspired by human-in-the-loop machine learning works [25, 22]).
  • ILSA updates only the Transformer: C1. Update the entire architecture; C2. Freeze the transformer and update other networks.
For the ablation study, one researcher runs five experiments for each model on the cereal pouring task. Each experiment consists of four trials with different task layouts. The task completion times of different models are shown in Fig. 4.
As shown in Fig. 4, our proposed method, ILSA, consistently decreases task completion times across trials. With just a few user interactions, ILSA lowers the average task completion time from 55.2 seconds in trial 1 to 35.6 seconds in trial 4, which is a 35.5% reduction. From trial 2 to trial 4, ILSA has the lowest average task completion time and one of the smallest standard deviations, demonstrating both effectiveness and repeatability.
In contrast, the method using uncorrected trajectories (A1) shows a smaller reduction in task completion time in trials 2 and 3, with larger standard deviations. This is mainly because user corrections for previous errors in the uncorrected trajectories are outliers that can cause overfitting. This can lead to errors in later trials, resulting in higher average task completion times and standard deviations.
Similarly, updating the model only on new data (B1), updating all neural networks in the architecture (C1), or freezing the transformer and updating other networks (C2) also leads to worse incremental learning and task performance. These ablated methods performed poorly on task layouts that are highly dissimilar to previously seen scenarios.
The three methods that use both new data and synthetic kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$ (B2, B3, B4) show a relatively steady decrease in average task completion times. However, compared to ILSA, they improve performance at a slower rate, indicating that they are not utilizing new data as efficiently as ILSA.
(a) Cereal pouring task
(b) Pill bottle storage task
IV-C User Study
We conducted a user study subject to a university-approved IRB protocol with 20 participants, none of whom reported having prior experience teleoperating robots. We test the following two hypotheses:
  • H2: ILSA improves task performance over time as a user repeatedly performs a task, in contrast to a static shared autonomy method.
Participants were divided into two groups. Group 1, consisting of 10 participants, used ILSA and pure teleoperation, while Group 2, consisting of 10 participants, used teleoperation and a static shared autonomy policy, which was pretrained on the initial synthetic kinematic trajectories and kept static across the four task trials.
IV-C1 Task Completion Times
IV-C2 User-reported Qualitative Metrics
In addition to capturing task completion time, we asked participants to report their perception of the robot control using 7-point Likert-type items (1 = strongly disagree; 4 = neutral; 7 = strongly agree) with items corresponding to “feel in control”, “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, “precise motions”, and “no errors”. Participants’ median responses at trial 4 are shown in Fig. 6.
For both tasks, most participants reported a higher value on most metrics for ILSA compared to teleoperation. By applying a Wilcoxon signed-rank test, for the cereal pouring task, we observe a statistically significant difference between ILSA and teleoperation on “low mental effort” ($p<0.01$), “intuitiveness” ($p<0.05$), “ease of task completion” ($p<0.01$), “smooth motions” ($p<0.01$) and “no errors” ($p<0.05$). For the pill bottle storage task, we observe a statistically significant difference between ILSA and teleoperation on all 7 metrics ($p<0.05$ for “feel in control” and “precise motions”, $p<0.01$ for “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, and “no errors”). These results support that users in general prefer ILSA over teleoperation, which provides evidence towards H1.
(a) Cereal pouring task
(b) Pill bottle storage task
Figure 6: Median Likert item responses for all methods after trial 4.
V CONCLUSIONS
In this paper, we introduce an innovative Incrementally Learned Shared Autonomy (ILSA) system that starts with pretraining on synthetic kinematic trajectories and then incrementally finetuned through user interactions. Our finetuning paradigm supports continuous improvement after each user interaction, balancing the integration of new data with the retention of prior knowledge. Through ablation and user studies, we compare ILSA to alternative designs, teleoperation, and a static shared autonomy method, demonstrating that ILSA is capable of enhancing task performance for complex multi-stage tasks over repeated task execution.

### v3

@@ III-C Robot Action Generation Model
Our robot action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$, shown in Fig. 2, produces two types of actions at each time step: an intermediate action $a_{m,t}=\mathcal{F^{\prime}}_{\theta^{\prime}}(s_{t})$ without user guidance and a final action $a_{f,t}=\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$ with user guidance. Here, $\mathcal{F^{\prime}}_{\theta^{\prime}}$ and ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$ represent different subsets of the overall model $\mathcal{F}_{\theta}$, as illustrated in Fig. 2.While only the action with user guidance is executed, having an intermediate output is crucial during our incremental learning process (Section III-D).
To pretrain $\mathcal{F}_{\theta}$, we use a set of simulated kinematic trajectories (Section III-B) and optimize a loss function designed to balance task completion and user intent alignment. This function includes terms ($\mathcal{L}_{demo,m}$, $\mathcal{L}_{demo,f}$) that align both intermediate and final actions with training data to ensure task completion, along with additional constraints ($\mathcal{L}_{direc}$, $\mathcal{L}_{order}$) to maintain consistency with user inputs. Since this model serves as the basis for ILSA’s assistive policy rather than its core technical contribution, which focuses on incremental adaptation to real-world challenges, we defer the full mathematical formulation of the loss functions and implementation details to Appendix A.
III-D Structured Incremental Learning Mechanism
To enable ILSA to adapt to unforeseen real-world challenges beyond pre-collected data, we introduce a structured incremental learning mechanism.
This process begins with how ILSA interacts with the user in real time: During deployment, at each time step, the user provides an action $a_{u,t}$ through the control interface, which is input into $\mathcal{F}_{\theta}$ to generate the final robot action $a_{f,t}$. Since real-world deployment may introduce unforeseen challenges beyond the scope of pretraining, the pretrained model may generate suboptimal actions when encountering these novel situations. To address this, we monitor the alignment between user inputs and model predictions by computing the cosine similarity between $a_{u,t}$ and $a_{f,t}$. If the similarity exceeds a threshold (empirically set to 0.5), $a_{f,t}$ is executed. Otherwise, we alert the user by briefly pausing the robot’s motion. If the user persists with the same input, the robot will execute $a_{u,t}$ instead of $a_{f,t}$. This process generates new interaction data, which we use to incrementally update $\mathcal{F}_{\theta}$.
While a straightforward approach would be to directly use these interaction trajectories for fine-tuning, this presents several issues: (1) suboptimal trajectories introduced by initial model limitations may be reinforced rather than corrected, (2) adapting to new scenarios risks forgetting of previously learned behaviors, and (3) learning from limited deployment data can lead to instability or reduce generalization ability. To address these challenges, we propose three key designs: Corrected Trajectory Supervision, Layered Supervision, and Partial Model Update, which together enable robust and efficient incremental learning.
III-D1 Corrected Trajectory Supervision
Our experiments show that when users overwrite $a_{f,t}$ with $a_{u,t}$, it is typically to correct suboptimal actions previously made by the system, which is usually caused by unforeseen task variations. For example, as illustrated in Fig. 1(a), the shelf above the cup is an unexpected obstacle that the model was not exposed to during training, leading the robot arm to collide with it while attempting to retrieve the cup. To recover, the user intervenes by commanding the robot to move downward, an action that deviates from the model’s expected behavior and requires overwriting $a_{f,t}$ with $a_{u,t}$. This correction is only necessary because the initial model, unaware of the obstacle, attempted an infeasible trajectory that goes up.
As illustrated in this example, raw trajectories contain both suboptimal policy actions and user corrections that wouldn’t be necessary without the suboptimal actions. Directly using these trajectories for training risks misleading the model into treating these corrective actions as optimal behavior to be replicated.
To address this, we generate a “corrected trajectory” for each raw trajectory observed during human use. As shown in Fig. 3, when the user overwrites $a_{f,t}$ (red segments in Fig. 3), let $t_{1}$ be the start of the overwrite and $t_{2}$ the time when the robot resumes executing $a_{f,t}$. Before $t_{1}$, if the robot executes $a_{f,t}$ for a period (blue segments), let $t_{0}$ be the start of that period. Let $p_{t_{0}}$, $p_{t_{1}}$, and $p_{t_{2}}$ represent the robot’s positions at these times. An example of these key positions is marked in Fig. 3.
III-D2 Layered Supervision
To ensure a smooth user experience, ILSA updates its policy after each interaction. However, frequent updates mean that each training step relies on only a small set of newly collected data, making the system susceptible to overfitting and catastrophic forgetting of previously acquired knowledge [9, 22, 15].
A naïve solution is to mix newly collected trajectories with the pretraining trajectories. However, this approach is problematic due to the severe data imbalance—especially in early updates when only a few new trajectories exist. A common alternative used in human-in-the-loop machine learning methods is to assign different weights to different types of data [21, 18, 17], but determining optimal weights is nontrivial and often task-specific.
Instead, we propose a layered supervision strategy that naturally integrates both new interaction data and simulated pretraining data without explicit reweighting:
  1. 1.
    Final Output Supervision: The new user-generated trajectories are used exclusively to supervise the final output action $a_{f,t}$, ensuring adaptation to real-world interactions.
  2. 2.
    Intermediate Output Supervision: Both simulated kinematic trajectories and new user-generated trajectories supervise the intermediate action $a_{m,t}$, maintaining the model’s generalization ability while still allowing adaptation.
This approach, illustrated by the red arrows in Fig. 2, ensures that ILSA remains adaptive while preserving knowledge from pretraining. Compared to explicit weighting schemes, layered supervision provides a more principled way to balance adaptation and generalization.
III-D3 Partial Model Update
To further enhance robustness, we update only a subset of $\mathcal{F}_{\theta}$ during fine-tuning. Specifically, we focus on the transformer component, which processes the task state $s_{t}$, while freezing other network modules. This design is motivated by the need for ILSA to efficiently adapt to previously unseen real-world variations—such as new obstacles or spatial constraints—without discarding valuable prior knowledge. Since the transformer extracts high-level task representations from $s_{t}$, fine-tuning it alone allows the policy to evolve with each user interaction while preserving generalization ability. Our ablation study (Section IV-B) demonstrates that this selective update strategy significantly outperforms full-model updates or modifying other components.
IV Experiments
IV-A Tasks and Experiment Settings
All experiments are conducted with a 6-DOF Kinova Gen3 robotic arm, and users use an Xbox controller to send commands to the robot.
As described in Section III-B, our pretraining phase relies on simulated kinematic trajectories that do not account for real-world complexities such as obstacles or spatial constraints. However, in real-world applications, previously unseen factors can fundamentally alter task execution. To evaluate ILSA’s ability to adapt in such scenarios, we design our experimental tasks to introduce obstacles and spatial constraints that significantly impact motion feasibility.
Specifically, we design two complex, multi-stage tasks to test the effectiveness of our model:
IV-A1 Cereal Pouring
This task introduces an obstacle (the shelf) that was not accounted for in pretraining.
IV-A2 Pill Bottle Storage
Open a drawer, pick up a pill bottle behind the drawer, and place it back inside.
In this task, the spatial constraints between the drawer and the pill bottle were not accounted for in pretraining, requiring ILSA to adapt its motion strategy to avoid unintended collisions.
IV-B Ablation Study
Before conducting the user study, we performed an ablation study to systematically evaluate ILSA’s structured fine-tuning mechanism. Specifically, we analyzed the impact of its three key components—Corrected Trajectory Supervision, Layered Supervision, and Partial Model Update. We evaluate the impact of these components through the following variations:
  • Corrected Trajectory Supervision: A. Fine-tune on original uncorrected trajectories.
  • Layered Supervision B1. Use new data only to supervise both $a_{f,t}$ and $a_{m,t}$; B2. Use both new data and pretraining data to supervise both $a_{f,t}$ and $a_{m,t}$; B3. Use new data and randomly selected 10 pretraining trajectories to supervise both $a_{f,t}$ and $a_{m,t}$; B4. Use both new data and pretraining data to supervise both $a_{f,t}$ and $a_{m,t}$, reweighted with $w_{new}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{new}|}$ and $w_{kinematic}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{kinematic}|}$, where $|D_{new}|$ and $|D_{kinematic}|$ represent the number of data points in the new and kinematic trajectories, respectively (This design is inspired by human-in-the-loop machine learning works [21, 18]).
  • Partial Model Update C1. Update the entire architecture; C2. Freeze the transformer and update other components.
For the ablation study, one researcher conducted five experiments for each model variant on the cereal pouring task. The tested designs were randomized and interleaved across experiments. To evaluate adaptation across diverse task layouts, object positions were randomized for each trial, resulting in a total of 20 unique task layouts across the five experiments. For fair comparison, all model variants were evaluated on the same sequence of object positions within each experiment. This ensures that performance differences arise from model adaptation rather than variations in task difficulty.
Across both the ablation and user studies (Section II-B), we primarily  [^1] Additional experimental results are available on our project website. evaluate task completion time as a direct and interpretable measure of user efficiency and system adaptability, which is particularly relevant in assistive settings where efficiency and ease of use directly impact user experience.
The task completion times of different models are shown in Fig. 4. As shown in the figure, our proposed method, ILSA, consistently decreases task completion times across trials. With just a few user interactions, ILSA lowers the average task completion time from 55.2 seconds in trial 1 to 35.6 seconds in trial 4, which is a 35.5% reduction, suggesting effective adaptation to unforeseen obstacles that were not accounted for during pretraining. From trial 2 to trial 4, ILSA has the lowest average task completion time and one of the smallest standard deviations, demonstrating both effectiveness and repeatability.
In contrast, the method using uncorrected trajectories (A1) shows a smaller reduction in task completion time in trials 2 and 3, with larger standard deviations. This is mainly because user corrections for previous suboptimal actions in the uncorrected trajectories are outliers that can cause overfitting. This can lead to errors in later trials, resulting in higher average task completion times and standard deviations.
Similarly, updating the model only on new data (B1), updating the entire architecture (C1), or freezing the transformer and updating other components (C2) also leads to worse incremental learning and task performance. These ablated methods performed poorly on task layouts that are highly dissimilar to previously seen scenarios.
The three methods that use both new data and simulated kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$ (B2, B3, B4) show a relatively steady decrease in average task completion times. However, compared to ILSA, they improve performance at a slower rate, indicating that they are not utilizing new data as efficiently as ILSA.
(a) Cereal pouring task
(b) Pill bottle storage task
IV-C User Study
To formally evaluate ILSA, We conducted a user study subject to a university-approved IRB protocol with 20 participants, none of whom reported having prior experience teleoperating robots. We test the following two hypotheses:
  • H2: ILSA effectively adapts to real-world challenges beyond the scope of training data, continuously improving task performance over time, in contrast to a static shared autonomy method.
Participants were divided into two groups. Group 1, consisting of 10 participants, used ILSA and pure teleoperation, while Group 2, consisting of 10 participants, used teleoperation and a static shared autonomy policy, which was pretrained on the initial simulated kinematic trajectories and kept static across the four task trials.
IV-C1 Task Completion Times
IV-C2 User-reported Qualitative Metrics
In addition to task completion time, we collected participants’ subjective evaluations using a custom-designed 7-point Likert-scale questionnaire (1 = strongly disagree; 4 = neutral; 7 = strongly agree) to assess their perception of robot control. The evaluated metrics included “feel in control”, “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, “precise motions”, and “no errors”. We note that these subjective evaluations focus on user experience and perceived control. Metrics like “precise motions” and “smooth motions” reflect how intuitive and seamless the system feels to users, rather than physical accuracy. This aligns with the primary goal of assistive systems: enhancing user experience and control. Given the multiple-trial, multiple-method nature of our study, we opted for this concise questionnaire instead of lengthier standardized assessments such as NASA-TLX to minimize participant burden and maintain engagement. The full questionnaire is available on our project website.
Participants’ median responses at trial 4 are shown in Fig. 6. For both tasks, most participants reported a higher value on most metrics for ILSA compared to teleoperation. By applying a Wilcoxon signed-rank test, for the cereal pouring task, we observe a statistically significant difference between ILSA and teleoperation on “low mental effort” ($p<0.01$), “intuitiveness” ($p<0.05$), “ease of task completion” ($p<0.01$), “smooth motions” ($p<0.01$) and “no errors” ($p<0.05$). For the pill bottle storage task, we observe a statistically significant difference between ILSA and teleoperation on all 7 metrics ($p<0.05$ for “feel in control” and “precise motions”, $p<0.01$ for “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, and “no errors”). These results support that users in general prefer ILSA over teleoperation, which provides evidence towards H1.
(a) Cereal pouring task
(b) Pill bottle storage task
Figure 6: Median Likert item responses for all methods after trial 4.
V CONCLUSIONS
This paper presents ILSA, an Incrementally Learned Shared Autonomy system that continuously refines its assistive policy through user interactions. A key contribution of ILSA is its structured fine-tuning mechanism, which enables per-interaction adaptation by efficiently leveraging limited new data while retaining prior knowledge, ensuring a stable balance between adaptability and generalization. Through ablation and user studies, we demonstrate that ILSA effectively handles unforeseen real-world challenges, including those that fundamentally alter task dynamics, such as unexpected obstacles or spatial constraints, and outperforms alternative designs, teleoperation, and static shared autonomy.
Appendix A Loss Function Details
Our loss function is defined as:
$\mathcal{L}=\alpha\mathcal{L}_{demo,m}+\beta\mathcal{L}_{demo,f}+\gamma\mathcal{L}_{direc}+\delta\mathcal{L}_{order}$ (1)
where $\alpha$, $\beta$, $\gamma$, and $\delta$ are weighting parameters empirically set to $\alpha=\beta=1$ and $\gamma=\delta=100$.
@@ Appendix A Loss Function Details
where $a_{r,t}$ is the robot action in the training trajectories. These two loss terms supervise the intermediate and final outputs respectively, guiding the robot to take actions towards task completion. To further ensure that the final executed action aligns with the user’s input, we introduce two additional loss terms, action direction loss $\mathcal{L}_{direc}$ and action ordering loss $\mathcal{L}_{order}$, to supervise the final output $a_{f,t}$.